Use a scale drawing to find the x- and y-components of the following

vectors. In each case, the magnitude of the vector and the angle,

measured counterclockwise, that it makes with the +x-axis are given.

a) magnitude 7.40 m, angle 300


; b) magnitude 15 km, angle 2250

; c)magnitude 9.30 cm, angle 3230


1
Expert's answer
2021-08-31T08:57:58-0400

a) magnitude 7.40 m, angle 300


 "a_x=7.4\cos 300^{\circ}=3.7\:{\rm m}\\
a_y=7.4\sin 300^{\circ}=-6.4\:{\rm m}"

b) magnitude 15 km, angle 2250=90


"a_x=15\cos 90^{\circ}=0\:{\rm km}\\
a_y=15\sin 90^{\circ}=15\:{\rm km}"

c)magnitude 9.30 cm, angle 3230=350


"a_x=9.3\cos 350^{\circ}=9.2\:{\rm cm}\\
a_y=9.3\sin 350^{\circ}=-1.6\:{\rm cm}"